An equilibrium mixture contains 0.300 mol of each of the products (carbon dioxide and hydrogen gas) and 0.200 mol of each of the reactants (carbon monoxide and water vapor) in a 1.00-L container.

CO+H20 <-> CO2 +H2

How many moles of carbon dioxide would have to be added at constant temperature and volume to increase the amount of carbon monoxide to 0.300 mol once equilibrium has been reestablished?
Answer #1

For this reaction, CO (g) + H2O (g) <------------> CO2 (g) + H2 (g)
Kc = [CO2] [H2] / [CO] [H2O] = 0.3 * 0.3 / 0.2*0.2 = 2.25

Kc = (0.2+x)*0.2 / 0.3*0.3 = 2.25
0.04 + 0.2x = 0.2025
0.2x = 0.2025 - 0.04 = 0.1625
x = 0.1625 / 0.2 = 0.8125
moles of CO2 added at constant temperature and volume = 0.8125 mol